if we assume that a given bus market is in perfect competition which charges a flat fare of $1 and if the formula for the total demand in the market is given by qd=250-60p where qd is the quantity demand at a given price. if we further assume constant returns to scale. what is the total market demand at the $1 flat fare?
1
Expert's answer
2015-04-09T08:38:02-0400
P = N$1, Qd = 250-60P
The total demand at the N$1 flat fare is: Qd = 250 - 60*1 = 190 thousands.
If the market is shared equally by 4 firms, the number of passengers per vehicle carried by each company is Q = 190/4 = 47.5 thousands.